Solve. Please show all of your work.
The houses on the north side of Polk Avenue are consecutive odd numbers. Bob and Laura are next-door neighbors and the sum of their house numbers is 524. Find both house numbers.

Answer by Maths68(1474)

Let
Bob's house number = x+1
Laura's house number = x+3
Sum of the numbers
x+1+x+3=524
2x+4=524
2x=524-4
2x=520
2x/2=520/2
x=260

Bob's house number = x+1 = 260+1 = 261
Laura's house number = x+3 = 260+3 = 263